# i-PI 3.0: a flexible, efficient framework for advanced atomistic simulations

### v1

@@ V.3 Melting of Charge Density Waves in 2D Materials
Downfolded lattice models aim to reduce the complexity of the full electronic problem by treating only a few low-energy electronic states (around the Fermi energy) explicitly and defining a simplified Hamiltonian where all other high-energy states are integrated out.[181] The parameters required for this dimensionality reduction can be obtained directly from ab initio calculations. There are several ways to include the effect of screening of the high-energy electrons on the several interaction terms related to the explicit degrees of freedom of the downfolded Hamiltonian,[182] and making appropriate choices in this respect is necessary to reproduce the fully ab initio potential energy surface of certain materials.[19]
@@
Figure 17: (a) Ratio of structure factors $S_{\text{CL}}/S_{\text{PI}}$ from classical and path integral MD performed on a $18\times 18$ supercell of monolayer 1H-TaS2. A value close to 1 (indicated in white) corresponds to minimal differences between classical and quantum simulations. The peaks at $\bm{q}=2/3\Gamma M$ and $\bm{q}=K$ for $T=50$ K are characteristic of the $3\times 3$ CDW phase. (d) Temperature-dependent structure factor $\langle S(\bm{q})\rangle_{T}$ at the characteristic CDW wavevector $\bm{q}=2/3\Gamma M$ obtained from classical MD replica-exchange MD (light blue) and path integral replica-exchange MD (blue). The experimental value (for the bulk) is represented by the vertical black dashed line. The shift of the inflection point of the PIMD curve toward the experimental value is due to nuclear quantum effects. Figure and caption adapted from Ref. 19, with permission.
In connection with i-PI, this capability allows the simulation of different types of molecular dynamics with such downfolded models. While machine-learning potentials could play a similar role, there is an undeniable added value of physical interpretability when employing these models. Moreover, even if downfolded models cannot be applied to all types of materials, for suitable systems they are straightforward to parametrize.
In Ref. [19], large-scale replica-exchange and path-integral replica exchange simulations were performed with the i-PI/elphmod combination in order to study the charge-density-wave (CDW) phase transition in monolayer 1H-TaS2. The downfolded model was obtained from PBE density-functional theory calculations. In particular, the phase-transition between the 3$\times$3 CDW and the symmetric structure was studied. We note that capturing the nuclear dynamics involved in such phase transitions calls for techniques that include non-perturbative anharmonic effects such as (PI)MD. In addition, as explicitly shown in Ref. [19], finite-size effects on the phase-transition are significant for small unit cells in these systems, calling for potentials that are efficient enough to allow the simulation of larger unit cells containing thousands of atoms.
Experimentally, the “melting” temperature of the 3 $\times$ 3 CDW in bulk TaS2 is measured at around 75 K.[183] While the phase-transition temperature is not known for the monolayer, it is accepted that it should not strongly deviate from the bulk temperature. In panels a and b of Fig. 17, we show the ratio between classical and quantum $\mathbf{q}$-resolved structure factors and the intensity of the CDW $3\times 3$ diffraction peaks with varying temperature, respectively. The inflection point of these curves is shown to correlate with a peak in heat capacity calculated from the simulations. Nuclear quantum effects decrease the transition temperature by around 20 K, bringing it in closer agreement with the experimental estimate.
Such simulations open the path for a closer analysis of particular phonon modes that contribute to the CDW phase transition, an analysis of the co-existence of different CDW phases at metastable regions, and the mapping of more complex CDW phase diagrams. Given the close connection between CDW phases and superconductivity, this is a promising field of study. An example on how to setup and run the elphmod code in connection with i-PI is included in the examples folders of the i-PI and the elphmod repositories.
@@ VI Conclusions
In this manuscript, we described the latest release of the i-PI code, a flexible and efficient framework for advanced atomistic simulations. We presented several new features implemented since the previous release, 5 years ago. Just as importantly, we discussed the efforts made by the core developers and contributors towards better code documentation and improved code stability and performance. In this regard, we showed extensive performance benchmarks demonstrating that i-PI’s current computational overhead is minimal when using state-of-the-art MLIPs for production simulations, and reported that the current version reduces between 2 and 10 times the overhead of several common setups. Considering also the availability of uncertainty quantification, that can be propagated from an ensemble of potentials to thermodynamic averages, i-PI is now an ideal companion for ML-powered atomistic simulations.
The ease by which new types of advanced simulations can be implemented, or even realized using the ability of i-PI to combine different forcefields and sampling algorithms, is demonstrated by the breadth of examples we showed in this paper, including simulations coupled to external fields, sampling of bosonic statistics and the estimation of spectroscopic observables by approximate quantum dynamics. We hope that this manuscript encourages more researchers to contribute their algorithms into i-PI, or simply to find in i-PI  a useful, reliable and friendly tool that facilitates incorporating advanced simulation techniques to realize complex and insightful simulations of matter.
@@ Acknowledgements.
Y.L. has been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) project number 467724959. V.K. acknowledges support from the Ernest Oppenheimer Early Career Fellowship and the Sydney Harvey Junior Research Fellowship, Churchill College, University of Cambridge. V.K is grateful for computational support from the Swiss National Supercomputing Centre under project s1209, the UK national high-performance computing service, ARCHER2, for which access was obtained via the UKCP consortium and the EPSRC grant ref EP/P022561/1, and the Cambridge Service for Data Driven Discovery (CSD3). T.B. acknowledges financial support from the Swiss National Science Foundation through the Early Postdoc Mobility Fellowship (grant number P2ELP2-199757). B.H. acknowledges support by the USA-Israel Binational Science Foundation (grant No. 2020083) and the Israel Science Foundation (grants No. 1037/22 and 1312/22). Y.F. was supported by Schmidt Science Fellows, in partnership with the Rhodes Trust. T.E.L. is supported by start-up funds from the University of Delaware Department of Physics and Astronomy. M.R. and E.S. acknowledge computer time from the Max Planck Computing and Data Facility (MPCDF), funding from the IMPRS-UFAST program and the Lise-Meitner Excellence program. M.R. thanks Jan Berges for a careful read of section V C. M.C., M.K. and D.T. acknowledge funding from the Swiss National Science Foundation (SNSF) under the projects CRSII5_202296 and 200020_214879. M.C. and D.T. also acknowledge the support from the MARVEL National Centre of Competence in Research (NCCR) M.C. acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme Grant No. 101001890-FIAMMA.
@@ VII Data Availability
The version of i-PI used in the examples and benchmarks can be obtained from https://github.com/i-pi/, using the tag v3.0.0-alpha2. The data required to reproduce the figures and the benchmark simulations are available at https://github.com/i-pi/ipiv3_data.

### v2

@@ V.3 Melting of Charge Density Waves in 2D Materials
Downfolded lattice models aim to reduce the complexity of the full electronic problem by treating only a few low-energy electronic states (around the Fermi energy) explicitly and defining a simplified Hamiltonian where all other high-energy states are integrated out.[200] The parameters required for this dimensionality reduction can be obtained directly from ab initio calculations. There are several ways to include the effect of screening of the high-energy electrons on the interaction terms related to the explicit degrees of freedom of the downfolded Hamiltonian,[201] and making appropriate choices in this respect is necessary to reproduce the fully ab initio potential energy surface of certain materials.[20]
@@
In connection with i-PI, this capability allows the simulation of different types of molecular dynamics with such downfolded models. While MLIPs could play a similar role, there is an added value of physical interpretability when employing these models. Moreover, even if downfolded models cannot be applied to all types of materials, for suitable systems they are straightforward to parametrize.
In Ref. 20, large-scale replica-exchange and path-integral replica exchange simulations were performed with the i-PI/elphmod combination in order to study the charge-density-wave (CDW) phase transition in monolayer 1H-TaS2. The downfolded model was obtained from PBE density-functional theory calculations. In particular, the phase transition between the 3$\times$3 CDW and the symmetric structure was studied. We note that capturing the nuclear dynamics involved in such phase transitions calls for techniques that include non-perturbative anharmonic effects such as (PI)MD. In addition, as explicitly shown in Ref. 20, finite-size effects on the phase transition are significant for small unit cells of these systems, calling for potentials that are efficient and accurate enough to allow the simulation of larger unit cells containing thousands of atoms.
Experimentally, the “melting” temperature of the 3 $\times$ 3 CDW in bulk TaS2 is measured at around 75 K.[202] While the phase-transition temperature is not known for the monolayer, it is accepted that it should not strongly deviate from the bulk temperature. In Fig. 17, we show the ratio between classical and quantum $\mathbf{q}$-resolved structure factors in panel (a) and the intensity of the CDW $3\times 3$ diffraction peaks with varying temperature in panel (b). The inflection point of these curves is shown to correlate with a peak in heat capacity calculated from the simulations. Nuclear quantum effects decrease the transition temperature by around 20 K, bringing it in closer agreement with the experimental estimate.
Such simulations open the path for a closer analysis of particular phonon modes that contribute to the CDW phase transition, an analysis of the co-existence of different CDW phases at metastable regions, and the mapping of more complex CDW phase diagrams. Given the close connection between CDW phases and superconductivity, this is a promising field of study. An example of how to set up and run the elphmod code in connection with i-PI is included in the examples/clients directory.
@@ VI Conclusions
In this paper, we described the current capabilities of the i-PI code. The fundamental reasoning behind the design decisions made during its early development, namely an architecture that allows rapid prototyping in Python, and utilities to facilitate implementing complicated sampling algorithms, still prove to be very useful. However, some of the underlying assumptions made in the past are no longer valid. For example, the widespread adoption of MLIPs has dramatically reduced the cost of energy evaluation relative to explicit electronic-structure calculations, making the overhead associated with the evolution of the atomic coordinates a more pressing concern. In this paper, we showed that relatively superficial changes in the implementation of some of the core classes in i-PI substantially accelerated the most severe bottlenecks, ensuring that i-PI can now be used in combination with state-of-the-art MLIPs without adding significant overhead, even when pushing the energy evaluation to high levels of parallelism.
At the same time, the modular design of i-PI makes it easy to implement radically new modeling ideas, bringing algorithmic speed-ups and new useful features to the codebase. A bosonic PIMD implementation with quadratic scaling, the implementation of the ring polymer instanton with explicit friction method, a framework for dynamics driven by an electric field, a framework for simulations of light-matter coupling, and a better interface with ensembles of MLIPs including uncertainty estimation, are the five examples we highlight. Many useful, and sophisticated, simulation workflows can also be realized without major implementation efforts, by simply combining existing methods with external scripts and suitable post-processing. We demonstrate this with three applications: path integral coarse-grained simulations for vibrational spectra, calculations of the 2D-IR spectrum of water, and the melting of 2D charge-density waves.
As MLIPs become faster and enable simulations with larger lengths and time scales, it might be necessary to revisit some of the deeper architectural decisions of i-PI to further reduce the overhead inherent to the serial Python-based implementation. The work that was made in preparation for this release demonstrates that these kinds of optimizations can coexist with a modular design, that supports the fast implementation of new and exciting methods for advanced molecular simulations.
@@ Acknowledgements.
Y.L. has been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) project number 467724959. V.K. acknowledges support from the Ernest Oppenheimer Early Career Fellowship and the Sydney Harvey Junior Research Fellowship, Churchill College, University of Cambridge. V.K is grateful for computational support from the Swiss National Supercomputing Centre under project s1209, the UK national high-performance computing service, ARCHER2, for which access was obtained via the UKCP consortium and the EPSRC grant ref EP/P022561/1, and the Cambridge Service for Data Driven Discovery (CSD3). T.B. acknowledges financial support from the Swiss National Science Foundation through the Early Postdoc Mobility Fellowship (grant number P2ELP2-199757). B.H. acknowledges support by the USA-Israel Binational Science Foundation (grant No. 2020083) and the Israel Science Foundation (grants No. 1037/22 and 1312/22). Y.F. was supported by Schmidt Science Fellows, in partnership with the Rhodes Trust. T.E.L. is supported by start-up funds from the University of Delaware Department of Physics and Astronomy. M.R. and E.S. acknowledge computer time from the Max Planck Computing and Data Facility (MPCDF), funding from the IMPRS-UFAST program and the Lise-Meitner Excellence program. M.R. thanks Jan Berges for a careful read of section V C. M.C., M.K. and D.T. acknowledge funding from the Swiss National Science Foundation (SNSF) under the projects CRSII5_202296 and 200020_214879. M.C. and D.T. also acknowledge the support from the MARVEL National Centre of Competence in Research (NCCR) M.C. acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme Grant No. 101001890-FIAMMA. We thank Eli Fields for helpful comments.
@@ VII Data Availability
i-PI can be obtained downloaded from https://github.com/i-pi/i-pi. The tag v3.0.0-beta was used to generate the examples and benchmarks reported in this manuscript. The data required to reproduce the figures and benchmark simulations are available at https://github.com/i-pi/ipiv3_data.